OpenStudy (anonymous):

A video store is selling old DVDs and video games. They charge $3 for each video game and $4 for each DVD. You buy a combination of 15 DVDs and games, and spend $51. How many of each did you buy? A. 9 games and 6 DVDs B. 10 games and 5 DVDs C. 11 games and 4 DVDs D. 12 games and 3 DVDs

OpenStudy (anonymous):

So you can solve the first equation for v to get v+d = 15 v = 15-d then plug this into the second equation to get 3v + 4d = 51 3(15 - d) + 4d = 51 45 - 3d + 4d = 51 d + 45 = 51 I'll let you finish up

OpenStudy (anonymous):

let v = # of video games bought d = # of dvds bought since you bought 15 items total, you know v+d = 15

OpenStudy (anonymous):

because "They charge $3 for each video game and $4 for each DVD" and you spent $51, you know that 3v + 4d = 51

OpenStudy (anonymous):

OK?

OpenStudy (anonymous):

Okay.. so it is c.

OpenStudy (anonymous):

Yep. Great job.